# Argumentative Debates for Transparent Bias Detection [Technical Report]

### v1

@@ 6 Arguing about Bias
Let $Q_{\mathcal{N}}\!=\!(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ be a local bias-QBAF for $g$ w.r.t. $\mathcal{N}$. Then the local bias-QBAF with critical questions for $g$ w.r.t. $\mathcal{N}$ is $\mathcal{Q}^{c}_{\mathcal{N}}\!=\!(\mathcal{A}^{\prime},{\mathcal{R}^{-}}^{\prime},\mathcal{R}^{+},\tau^{\prime})$ with $\mathcal{A}^{\prime}\!=\!\mathcal{A}\!\cup\!CQ$, ${\mathcal{R}^{-}}^{\prime}\!=\!\mathcal{R}^{-}\!\cup\!\{(a,\mathrm{Adv}_{g}),(a,\mathrm{Disadv}_{g})\!\mid\!a\!\in\!CQ\}$ and $\tau^{\prime}$ is such that for each $a\not\in CQ$, $\tau^{\prime}(a)=\tau(a)$ and for each $a\in CQ$, $\tau^{\prime}=f^{a}$ for functions $f^{a}$ as follows:
  • $f^{s}\!:\!\mathbb{N}\!\rightarrow\![0,1]$ takes in input $|\mathcal{N}|$ and: 1. it is monotonically decreasing, 2. $f^{s}(1)\!=\!1$, and 3. $\lim_{x\!\rightarrow\!\alpha}f^{s}(x)\!=\!0$ for some significance threshold $\alpha\!\in\!\mathbb{N}\!\cup\!\{\infty\}$;
  • for $S\!\subseteq\!\mathcal{D}$, $f^{o}\!:\!2^{\mathcal{D}}\!\rightarrow\!\{0,1\}$ takes in input $\mathcal{N}$ and is such that $f^{o}(\mathcal{N})\!=\!1$ if $\mathcal{N}$ is $S$-objective, and $f^{o}(\mathcal{N})\!=\!0$ else;
  • $f^{d}\!:\![0,1]\!\rightarrow\![0,1]$ takes in input the entropy $H_{\mathcal{N}}(g)$ of $\mathcal{N}$ and: 1. it is monotonically decreasing, 2. $f^{d}(0)\!=\!1$ and 3. $f^{d}(\beta)\!=\!0$ for some entropy threshold $\beta\in(0,1]$.
@@
Similarly to Section 6, properties of $\sigma$ guarantee that the new arguments behave as intended. We do not go into formal detail but observe that, if $\operatorname{agg}$ and $\operatorname{infl}$ satisfy 1. balance, then $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will remain unaffected by critical questions that do not apply (base score 0), and 2. monotonicity, then the strength of $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will decrease monotonically w.r.t. the strength (i.e. base score) of $s,o,d$.
@@
We combine local bias-QBAFs with critical questions across neighbourhoods $\mathcal{N}_{1},\dots,\mathcal{N}_{m}$, using the argument scheme in Figure 3, as follows. The conclusion yields a single argument ($\mathrm{bias}_{g}$}). We make the following assumptions: 1. we have no prior information about biases of the classifier and 2. in the absence of information, we see the classifier as unbiased (in dubio per reo). We capture this by setting the base score of $\mathrm{bias}_{g}$} to 0, meaning that we have no reason to believe there is a bias against $X_{p}\!=\!g$ to start with. The minor premise formulates that $X_{p}\!=\!g$ is disadvantaged in the neighbourhoods. We instantiate the detected disadvantage with two arguments per neighbourhood ($\mathrm{Adv}_{g}^{i}$ and $\mathrm{Disadv}_{g}^{i}$ in $\mathcal{N}_{i}$), giving $2m$ arguments. We use the major premise to identify an attack/support from each $\mathrm{Adv}_{g}^{i}$$\mathrm{Disadv}_{g}^{i}$ against $\mathrm{bias}_{g}$}. Formally:
@@
Again, properties of $\sigma$ can guarantee that local bias-QBAFs affect $\mathrm{bias}_{g}$ as expected. In particular, $\mathrm{bias}_{g}$ will remain unaffected by fully rejected local bias arguments (strength 0) and its strength will increase/decrease monotonically w.r.t. local $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ arguments. Formally:
@@ 7 Experiments
We conduct experiments with: 1. synthetically biased classifiers (two hand-crafted and one resulting from training on the Adult Census Income (ACI) dataset (Becker and Kohavi 1996), also used for testing ); 2. trained classifiers (logistic regression, but our method applies to any classifier); 3. LLMs (ChatGPT-4o). For training (2) and for testing (2,3) we use the Bank Marketing (Moro, Rita, and Cortez 2014) and COMPAS (ProPublica 2016; Angwin et al. 2016a) datasets. The datasets we choose are commonly used in the fairness literature (Räz 2024; Fawkes et al. 2024; Ehyaei, Farnadi, and Samadi 2024). We consider also LLMs as they are prone to societal biases due to their exposure to human-generated data (Ma et al. 2025; Xiang 2024). In all experiments, we compare with the argumentative approach by (Waller, Rodrigues, and Cocarascu 2024) (IRB in short). We report results with quadratic energy (DF-QuAD gave similar results). We report results for 1000 randomly sampled individuals and their KNN-neighbourhoods (Cover and Hart 1967), drawn from the test sets of the chosen datasets.
@@
We developed three synthetic models to obtain an objective ground truth in the setting of the ACI dataset: two globally biased models (Global 1 and Global 2) and one locally biased model (Local 1). Global 1 and 2 are hand-crafted decision trees (see Figures 5, 6 in Appendix C): the former predicts a negative label for all female individuals and a positive label for all others; the latter predicts a negative label for all black female individuals. For Local 1, we trained a logistic regression model on the ACI training set. In contrast to the global models, Local 1 operates at a neighbourhood level, changing the prediction to the negative label for female individuals if their neighbourhood exhibits an $\epsilon$-bias against $gender=female$ in the originally trained model (see Figure 7 in Appendix C).
@@
Bias Across Neighbourhoods. We consider multipleneighbourhoods of varying sizes (50, 100, 200). [^3] We run experiments with other sizes, ranging from $K=10$ to $K=200$ in intervals of 10. See Appendix C.2. Table 2 shows that our method achieves perfect performance for all classifiers. IRB fails for Global 2 for the same reasons as before.
@@
Table 3(a) gives results drawn from COMPAS, showing that our method identifies substantially more biased cases against African-American individuals (77 vs. 2), in line with the literature (Angwin et al. 2016b; Khademi and Honavar 2020), while detecting fewer or no biased cases for the other groups compared to IRB.Table 3(b) gives results for individuals identified as biased against specific age and marital status groups in the Bank Marketing dataset. In particular, for “YoungOrOld” age (younger than 25 years of age or older than 60 years of age), “Married”, and “Single” marital features, our method detects more biased cases than IRB. We show examples in Appendix C.
@@ 8 Argumentative Debates
Figure 4: Example of an Argumentative Debate (the debate in the yellow/blue boxes stems from the ones in Figure 1).
@@ 9 Conclusion
We presented a novel model-agnostic approach for bias detection for analysing both local and global biases through argument schemes and quantitative bipolar argumentation. Our approach addresses a critical need for algorithmic fairness, by integrating transparency into bias detection, while also exhibiting desirable properties, linked to established properties for quantitative bipolar argumentation, and performance advantages over an argumentative baseline.
We discussed how our approach may lead to argumentative debates: future work will focus on the realisation thereof within and across agents, as well as between agents and humans, including with user studies.

### v2

@@ 6 Arguing about Bias with ABIDE
Let $Q_{\mathcal{N}}\!=\!(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ be a local bias-QBAF. The corresponding local bias-QBAF with critical questions is $\mathcal{Q}^{c}_{\mathcal{N}}\!=\!(\mathcal{A}^{\prime},{\mathcal{R}^{-}}^{\prime},\mathcal{R}^{+},\tau^{\prime})$ with $\mathcal{A}^{\prime}\!=\!\mathcal{A}\!\cup\!CQ$, ${\mathcal{R}^{-}}^{\prime}\!=\!\mathcal{R}^{-}\!\cup\!\{(a,\mathrm{Adv}_{g}),(a,\mathrm{Disadv}_{g})\!\mid\!a\!\in\!CQ\}$ and $\tau^{\prime}$ is such that $\tau^{\prime}(a)=\tau(a)$ for all $a\not\in CQ$, and
  • $\tau^{\prime}(s)=f^{s}(|\mathcal{N}|)$ for a monotonically decreasing function $f^{s}\!:\!\mathbb{N}\!\rightarrow\![0,1]$ such that $f^{s}(1)\!=\!1$, and $\lim_{x\!\rightarrow\!\alpha}f^{s}(x)\!=\!0$ for some significance threshold $\alpha\!\in\!\mathbb{N}\!\cup\!\{\infty\}$;
  • $\tau^{\prime}(o)=f^{o}(\mathcal{N})$ where $f^{o}(\mathcal{N})\!=\!0$ if $\mathcal{N}$ is $S$-objective, and $f^{o}(\mathcal{N})\!=\!\gamma$ else, where $\gamma\in(0,1]$ is an objectivity weight;
  • $\tau^{\prime}(d)=f^{d}(H_{\mathcal{N}}(g))$ where for $f^{d}\!:\![0,1]\!\rightarrow\![0,1]$ is monotonically decreasing and satisfies $f^{d}(0)\!=\!1$ and $f^{d}(\beta)\!=\!0$ for some entropy threshold $\beta\in(0,1]$.
@@
Similarly to Section 6, properties of $\sigma$ guarantee that the new arguments behave as intended. We do not go into formal detail but observe that, if $\operatorname{agg}$ and $\operatorname{infl}$ satisfy balance, then $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will remain unaffected by critical questions that do not apply (base score 0), and, if they satisfy monotonicity, then the strength of $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will decrease monotonically w.r.t. the strength (i.e. base score) of $s,o,d$.
@@
Finally, we combine multiple local bias-QBAFs for different neighbourhoods $\mathcal{N}_{1},\dots,\mathcal{N}_{m}$, as per Figure 3. Their individual $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ arguments attack/support a global bias argument $\mathrm{bias}_{g}$}. We make the following assumptions: we have no prior information about biases of the classifier and, in the absence of information, we take that the classifier is unbiased. We capture this by setting the base score of $\mathrm{bias}_{g}$} to 0.
@@
Again, properties of $\sigma$ can guarantee that local bias-QBAFs affect $\mathrm{bias}_{g}$ as expected. In particular, $\mathrm{bias}_{g}$ will remain unaffected by fully rejected local bias arguments (strength 0) and its strength will increase/decrease monotonically w.r.t. local $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ arguments.
@@ 7 Experiments
We conduct experiments with: 1. synthetically biased classifiers (two hand-crafted and one resulting from training on the Adult Census Income (ACI) dataset (Becker and Kohavi 1996), also used for testing ); 2. trained classifiers (logistic regression, but our method applies to any classifier); 3. LLMs (ChatGPT-4o). For training (2) and for testing (2,3) we use the Bank Marketing (Moro, Rita, and Cortez 2014) and COMPAS (ProPublica 2016; Angwin et al. 2016) datasets. The datasets we choose are commonly used in the fairness literature (Räz 2024; Fawkes et al. 2024; Ehyaei, Farnadi, and Samadi 2024). We consider also LLMs as they are prone to societal biases due to their exposure to human-generated data (Ma et al. 2025; Xiang 2024). In all experiments, we compare ABIDE with the argumentative approach by (Waller, Rodrigues, and Cocarascu 2024) (IRB in short). We report results with quadratic energy (DF-QuAD gave similar results). We report results for 1000 randomly sampled individuals and their KNN-neighbourhoods (Cover and Hart 1967), drawn from the test sets of the chosen datasets. A sensitivity analysis evaluating the model’s performance across varying parameters is provided in (Ayoobi et al. 2025a).
@@
We developed three synthetic models to obtain an objective ground truth in the setting of the ACI dataset: two globally biased models (Global 1 and Global 2) and one locally biased model (Local 1). Global 1 and 2 are hand-crafted decision trees (see Figures 5, 6 in Appendix 3): the former predicts a negative label for all female individuals and a positive label for all others; the latter predicts a negative label for all black female individuals. For Local 1, we trained a logistic regression model on the ACI training set. In contrast to the global models, Local 1 operates at a neighbourhood level, changing the prediction to the negative label for female individuals if their neighbourhood exhibits an $\epsilon$-bias against $gender=female$ in the originally trained model (see Figure 7 in (Ayoobi et al. 2025a)).
@@
Bias Across Neighbourhoods. We consider multiple neighbourhoods of varying sizes (50, 100, 200). [^3] We run experiments with other sizes, ranging from $K=10$ to $K=200$ in intervals of 10. See (Ayoobi et al. 2025a) for details. Table 2 shows that our method achieves perfect performance for all classifiers. IRB fails for Global 2 for the same reasons as before.
@@
Table 3(a) gives results drawn from COMPAS, showing that our method identifies substantially more biased cases against African-American individuals (77 vs. 2), in line with the literature (Angwin et al. 2016; Khademi and Honavar 2020), while detecting fewer or no biased cases for the other groups compared to IRB. Table 3(b) gives results for individuals identified as biased against specific age and marital status groups in the Bank Marketing dataset. In particular, for “YoungOrOld” age (younger than 25 years of age or older than 60 years of age), “Married”, and “Single” marital features, our method detects more biased cases than IRB. We show examples in (Ayoobi et al. 2025a).
@@ 8 Argumentative Debates
Figure 4: Example of an Argumentative Debate (the debates in the yellow/blue boxes stem from the corresponding boxes in Figure 1).
@@ 9 Conclusion
We presented ABIDE, a novel model-agnostic approach for bias detection for analysing both local and global biases through argument schemes and quantitative bipolar argumentation. Our approach addresses a critical need for algorithmic fairness, by integrating transparency into bias detection, while also exhibiting desirable properties, linked to established properties for quantitative bipolar argumentation, and performance advantages over an argumentative baseline.
We discussed how ABIDE can support debates about bias. Future work will focus on the realisation thereof within and across agents, as well as between agents and humans, including with user studies. It will be especially interesting to explore how humans can contribute to the debates, specifically by contesting existing arguments and adding new arguments to the debate, in the spirit of (Leofante et al. 2024).
Acknowledgements
Ayoobi, Rapberger, and Toni were funded by the ERC under the EU’s Horizon 2020 research and innovation programme (ADIX, grant number 101020934). Toni was also funded by J.P. Morgan and by the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme.